How to create a quick graph and solve using elimination or substitution.
3x-10y=16
x=-6y

Answer by meetkisai(17)

# by substitution
given, x = -6y
hence,
3(-6y) - 10y = 16
-18y - 10y = 16
-28y = 16
=> y=-16/28
=> y= -4/7 and x = 24/7